OpenStudy (anonymous):

True or False Given f(x)=x-4, there is a zero of f(x) at x=0

OpenStudy (zepp):

False, when x = 0, you are finding the y intercept, zeroes are x intercepts

jimthompson5910 (jim_thompson5910):

When you plug in x = 0, you get f(x) = x-4 f(0) = 0-4 f(0) = -4 Since this result is NOT zero, this means that x = 0 is NOT a zero. So the statement is false.

OpenStudy (zepp):

Here's the graph http://puu.sh/r0ad

OpenStudy (anonymous):

Thanks everyone :)